OpenStudy (anonymous):

Can someone tell me if I got these right?? Question 4 (Multiple Choice Worth 1 points) Simplify completely quantity 8 x minus 56 over quantity x squared minus 49 divided by quantity x minus 6 over quantity x squared plus 11 x plus 28 4 open parentheses x plus 4 close parentheses over quantity x minus 3 8 open parentheses x plus 2 close parentheses over quantity x minus 3 8 open parentheses x plus 4 close parentheses over quantity x minus 7 8 open parentheses x plus 4 close parentheses over quantity x minus 6 I picked A.

OpenStudy (anonymous):

Simplify completely quantity 6 x squared minus 54 plus 84 over quantity 8 x squared minus 40 x plus 48 divided by quantity x squared plus x minus 56 over quantity 2 x squared plus 12 x minus 32 2 open parentheses x minus 3 close parentheses over 3 open parentheses x minus 2 close parentheses 2 open parentheses x minus 2 close parentheses over 3 open parentheses x minus 3 close parentheses 3 open parentheses x minus 3 close parentheses over 2 open parentheses x minus 2 close parentheses 3 open parentheses x minus 2 close parentheses over 2 open parentheses x minus 3 close parentheses I picked C

OpenStudy (anonymous):

Simplify completely quantity x squared minus 4 x plus 4 over quantity x squared plus 10 x plus 25 times quantity x plus 5 over quantity x squared plus 3 x minus 10 quantity x minus 2 over quantity x plus 5 x minus 2 all over the quantity x plus 5 end quantity squared the quantity x minus 2end quantity squared all over x plus 5 I picked B

OpenStudy (luigi0210):

Posting one at a time will make people more likely to help.

OpenStudy (anonymous):

Sorry, oh.

OpenStudy (anonymous):

how?

OpenStudy (luigi0210):

@Jhannybean where are you smartypants? >.<

OpenStudy (jhannybean):

the "@ " symbol

OpenStudy (anonymous):

oh..

OpenStudy (jhannybean):

Ex: @Luigi0210

OpenStudy (anonymous):

I seee now,

OpenStudy (luigi0210):

Yes?

OpenStudy (anonymous):

?

OpenStudy (jhannybean):

lets start with the first question

OpenStudy (anonymous):

okay

OpenStudy (luigi0210):

Well, @dreamer05 you're in good hands now. Time to hit the old trail. Good luck!

OpenStudy (anonymous):

thanks

OpenStudy (jhannybean):

Simplify completely quantity 8 x minus 56 over quantity x squared minus 49 divided by quantity x minus 6 over quantity x squared plus 11 x plus 28 (8x - 56)/ (x^2 -49) / (x-6)/(x^2 +11x +28) We can rewrite this as \[\large \frac{8x-56}{x^2-49} \cdot \frac{x^2+11x+28}{x-6}\]

OpenStudy (anonymous):

Ok.

OpenStudy (luigi0210):

Now factor and simplify!

OpenStudy (anonymous):

I did, that's why I am not sure if I got them right.. I'm not so good when it comes to fractions, that's why I wanna see if I got them right, and if I didn't i wanted it explained to me.

OpenStudy (jhannybean):

\[\large \frac{8\color{green}{(x-7)}}{\large \color{green}{(x-7)}\color{red}{(x+7)}}\cdot \frac{\color{red}{(x+7)}(x+4)}{x-6}\]

OpenStudy (jhannybean):

Now what can we cancel out here?

OpenStudy (anonymous):

the red? so the x+7?

OpenStudy (luigi0210):

When Jhanny's here.. an explanation is unavoidable >.>

OpenStudy (jhannybean):

And what else? :)

OpenStudy (jhannybean):

Youre on the right track.

OpenStudy (anonymous):

& the x-7

OpenStudy (jhannybean):

good job. \[\large \frac{8\color{green}{\cancel{(x-7)}}}{\large \color{green}{\cancel{(x-7)}}\color{red}{\cancel{(x+7)}}}\cdot \frac{\color{red}{\cancel{(x+7)}}(x+4)}{x-6}\]

OpenStudy (anonymous):

so its 8 x (x+4) over x-6

OpenStudy (jhannybean):

Re simplify by taking out all the stuff we've canceled. What are you left with?

OpenStudy (jhannybean):

Yay! :)

OpenStudy (anonymous):

so thats one?

OpenStudy (jhannybean):

a. 4 open parentheses x plus 4 close parentheses over quantity x minus 3 b. 8 open parentheses x plus 2 close parentheses over quantity x minus 3 c. 8 open parentheses x plus 4 close parentheses over quantity x minus 7 d. 8 open parentheses x plus 4 close parentheses over quantity x minus 6

OpenStudy (jhannybean):

So which one seems like the right choice?

OpenStudy (anonymous):

The last one

OpenStudy (jhannybean):

Good job.

OpenStudy (anonymous):

(:

OpenStudy (jhannybean):

Can you try the next one?

OpenStudy (anonymous):

I'll try.

OpenStudy (jhannybean):

break down the sentence into parts.Ask yourself, "how would i write 6 x squared minus 54 plus 84"?

OpenStudy (anonymous):

combine like terms?

OpenStudy (jhannybean):

Mmhmm, Most definitely.

OpenStudy (anonymous):

14x -94+132? I don't think that's right.

OpenStudy (jhannybean):

I'll write up the problem for you, the rest of it will be up to you! Just follow the same techniques we used in the last problem, alrighty? :)

OpenStudy (anonymous):

Okay c:

OpenStudy (jhannybean):

6 x squared minus 54 plus 84 over quantity 8 x squared minus 40 x plus 48 divided by quantity x squared plus x minus 56 over quantity 2 x squared plus 12 x minus 32 \[\large \frac{6x^2 -54 +84}{8x^2 -40x + 48}\cdot \frac{2x^2+12x-32}{x^2+x-56}\]

OpenStudy (anonymous):

8x^2 -54+12x+52 ----------------- 9x^2-39x-8

OpenStudy (anonymous):

17x^2-27x-10

OpenStudy (anonymous):

So is it A?

OpenStudy (anonymous):

@Jhannybean ?

OpenStudy (jhannybean):

\[\large \frac{6x^2 -54 +84}{8x^2 -40x + 48}\cdot \frac{2x^2+12x-32}{x^2+x-56}\]\[\large\frac{6(x^2-9x+16)}{8(x^2-5x+6)}\cdot \frac{2(x^2+6x-16)}{(x+8)(x-7)}\] Hope you can do the rest! I'm headed off OS now. Goodnight!

OpenStudy (anonymous):

Mkai night.

OpenStudy (luigi0210):

Goodnight Jhanny!

OpenStudy (luigi0210):

@dreamer05 did you finish the question?

OpenStudy (anonymous):

Yes, I only got that one wrong. I was able to get the rest right. Thanks. (:

OpenStudy (luigi0210):

So all your questions got answer?

OpenStudy (anonymous):

Yes they did. Thank you.

OpenStudy (luigi0210):

Awesome! And pfft, thanks Jhanny not me. :P

OpenStudy (anonymous):

(:

OpenStudy (luigi0210):

If you ever need anything just call us(Jhanny) up! :D Have a good rest of the day. :)

OpenStudy (anonymous):

You mean night, it's bed time for me lol but you too dear.